the sum of the digits of a two digit number is 8. When the number with the same digits reversed is subtracted from the number,the answer is 18. Find the number
Answer by lwsshak3(11628)

the sum of the digits of a two digit number is 8. When the number with the same digits reversed is subtracted from the number,the answer is 18. Find the number
**
let t=ten's digit of original number
let u=unit digit of original number
t+u=8
u=8-t
..
10t+u-(10u+t)=18
10t+8-t-(10(8-t)+t)=18
10t+8-t-(80-10t+t)=18
10t+8-t-80+10t-t=18
18t=90
t=5
u=8-t=3
ans:
original number=10t+u=50+3=53